46.3 mL of a 0.0999 M solution of this acid (HOOC-(CH2)2-COOH) is titrated to the equivalence point with 0.15 M KOH. Write the balanced equation and determine the volume of the base needed to completely neutralize the acid.
Answer #1

C 4H6O4(aq) + 2KOH(aq)   C 4H4O4Na2(aq) + 2H 2O(l)

Using, M1V1 (acid) = M2V2 (base)

0.0999×46.3 = 0.15×V2

V2 = 0.0999×46.3/0.15 = 30.84 ml

Please give feedback!